solve the quadratic equation by completing the square
x^2 - 6x - 3=0

Answer by malakumar_kos@yahoo.com(315)


solve the quadratic equation by completing the square
x^2 - 6x - 3=0

x^2-6x-3 = 0 reduce it to the std form i.e a^2+2ab+b^2 = (a+b)^2
x^2-2.3x = 3 compare with the std form we get 2ab = 2.3x so a= x & b = 3
add (b)^2 = (3)^2 on both sides
x^2- 2.3x+(3)^2 = 3+(3)^2
(x-3)^2 = 3+9
(x-3)^2 = 12
taking sq.rt on both sides
(x-3) = +-sq.rt 12
x = +-2 sq.rt3 +3
required solution is x = 3+-2 (sq.rt 3)